What volume of 0.259 M H2SO4 is needed to react with 47.1 mL of 0.116 M NaOH? The equation is

H2SO4(aq)+2NaOH(aq)→Na2SO4(aq)+2H2O(l)
Answer #1

using the dilution law M1V1 / n1 = M2V2 / n2

here M1& M2 are molarities

V1&V2 are volumes

n1 & n2 are no of moles

H2SO4 + 2 NaOH ------------------> Na2SO4 + 2 H2O

for H2SO4 => M1 - 0.259 M , V1 - ? , n1 - 1

for NaOH => M2 - 0.116 M , V2 - 47.1mL , n2 - 2

M1V1 / n1 = M2V2 / n2

V1 = M2V2n1 / n2M1

V1 = 0.116 * 0.0471 L * 1 / 2*0.259

V1 = 0.010547 L

V1 = 10.55 mL

answer => 10.55 mL